Marty leaves the airport in his private plane and flies due east at 186 mph. Two hours later, a jet leaves the same airport and flies due east at 434 mph. When will the jet overtake Marty's plane?
Answer by checkley71(8403)

186T=434(T-2)
186T=434T-868
186T-434T=-868
-248T=-868
T=-868/-248
T=3.5 HOURS AFTER THE FIRST PLANE LEAVES THEY MEET.